OpenStudy (anonymous):

If csc x = 7/3 and cot x = 2sqrt(10)/3, find sec x

TheSmartOne (thesmartone):

HI there :)

OpenStudy (anonymous):

Hi there again. Got up to another problem and got confused.

TheSmartOne (thesmartone):

So.. one sec.

TheSmartOne (thesmartone):

sin(\(\theta\)) = Opposite / Hypotenuse cos(\(\theta\)) = Adjacent / Hypotenuse tan(\(\theta\)) = Opposite / Adjacent

OpenStudy (anonymous):

So I would need to find the reciprocal of csc and cot?

OpenStudy (anonymous):

If think the answer is sec x = 7sqrt(10)/20

TheSmartOne (thesmartone):

And csc(\(\theta\)) = Hypotenuse / Opposite sec(\(\theta\)) = Hypotenuse / Adjacent cot(\(\theta\)) = Adjacent / Opposite

TheSmartOne (thesmartone):

Sorry OS lagged and it took forever to post that -_-

TheSmartOne (thesmartone):

So for csc x = 7/3 that means Hypotenuse is 7 and Opposite is 3.

TheSmartOne (thesmartone):



TheSmartOne (thesmartone):

If we solve for x, we get the adjacent side. And we already know the hypotenuse